OpenStudy (anonymous):

find the average rate from 3 to 5 for the following functions: 9t^2 anyone know it?

OpenStudy (anonymous):

Average rate = use integrals: \[1/(B-A) * \int\limits_{A}^{B} f(x) dx\] B = 5 A = 3\[1/2 * \int\limits_{3}^{5}9t^2 dt = .5*3t^3|_{3}^5 = 147\]

OpenStudy (anonymous):

i never heard of that lol im in pre cal dude

OpenStudy (anonymous):

Okay, so you could just average the values:\[f(t) = 9t^5\]\[Rate_{average} = f(5)-f(3)/2 = 225-81/2 = 72\]

OpenStudy (anonymous):

isnt it t^2

OpenStudy (anonymous):

That's what I meant. I did the calculations right, I just put 9t^5 :D

OpenStudy (anonymous):

kk lol